OpenStudy (anonymous):

are u = <1, -2>, v = <-4, 8> parallel?

OpenStudy (zehanz):

two vectors u = <u1, u2> and v = <v1, v2> are parallel if there is a number k such that v=u*k. This means: v1=k*u1 and v2=k*u2. This is easy to check with your u and v.